OpenStudy (anonymous):

Simplify this expression. 12x^2(3x^4 - 2x)

OpenStudy (anonymous):

\[12x^2(3x^4-2x)=36x^6-24x^3=6x^3(6x^3-4)\]

OpenStudy (anonymous):

Continuing from what @ConradBarczyk said, you can actually factor out a 12x^3. 36x^6−24x^3 = 12x^3(3x^3 - 2)

OpenStudy (anonymous):

ahh yes i knew there was a bigger factor i rushed :)

OpenStudy (anonymous):

lol. Yeah :)